# Exploring the Non-Markovian Dynamics in Depolarizing Maps

### v1

@@ Abstract
The non-Markovian depolarizing channel is explored from the perspective of understanding its non-Markovian behaviour as well as the occurrence of singularities. The study brings together the various ways to identify and quantify non-Markovianity. This includes dynamical techniques such as quantum information backflow witness, BLP, RHP and HCLA measures. In addition, geometrical visualization of non-Markovian effects is presented using the variation in the volume of accessible states during dynamical evolution. Further, a trajectory-based visualization of the dynamical map, of the depolarizing channel, within the parameter space is presented. The trajectories traced during evolution, demonstrate the loss of CP-divisibility and the emergence of non-Markovianity under systematic variations of the system parameters.
@@ I Introduction
The dynamical evolution of the open quantum system [1, 2], a composite system consisting of a smaller sub-system of interest and an interacting larger sub-system called environment, is of interest to the field of quantum information [3, 4] due to arising memory effects and non-Markovianity [5, 6, 7, 8, 9, 10, 11, 12]. The concept of the completely positive trace preserving (CPTP) map is used to study such evolution, which maps the system’s initial state $\rho_{S}(0)$ to the evolved final state $\rho_{S}(t)$ [3, 13]
@@
Alternately, CPTP maps are called quantum channels and can be described using operator sum representation (Kraus formalism) [14, 15, 16, 17]. The mapped evolution by the CPTP map can be classified as Markovian or non-Markovian. The Markovian evolution implies that system evolution is independent of past events and is dependent on only the present. On the contrary, such independence is violated, or the environment is found to pertain to the history of system evolution for non-Markovian evolution.
To identify non-Markovianity, traditionally, one resorts to the loss of CP-divisibility [7] or non-monotonic increase in the distinguishability of a pair of states during evolution [6]. Even CP-divisible processes can exhibit non-Markovain effects [9, 10]. The increase in distinguishability is indicative of information backflow from environment to system, indicating non-Markovianity. The Breuer-Laine-Piilo (BLP) measure [5] of non-Markovianity uses distinguishability to quantify non-Markovianity. CP-divisibility allows dynamics to be studied in discrete time intervals; the inability to do so is characteristic of a CP-indivisible or non-Markovan process. The dynamical map is said to be CP-indivisible, if the intermediate map, mapping intermediate evolution, is not completely positive (NCP), possessing at least one negative eigenvalue for the corresponding Choi matrix. The intermediate NCP map is indicative of system-environment correlation. The Rivas-Hulga-Plenio (RHP) measure [18, 7] uses the previously mentioned criteria to quantify non-Markovianity. Similarly, the Hall-Creser-Li-Anderson (HCLA) measure [19] quantifies non-Markovian effects but relies on negative time-dependent decay rates, which are indicative of non-Markovianity. Both measures are equivalent to each other and provide similar insights into non-Markovian processes.
@@
Our goal in this article is to answer the following two questions: (1) Under what conditions the depolarizing map becomes non-Markovian?, (2) Can we quantify the amount of this non-Markovianity?
@@
In order to answer these questions, both dynamical, based on CP-divisibility, distinguishability, negative time-dependent decay rates, quantum information backflow witness, and geometrical techniques are employed, and a comprehensive understanding of the non-Markovianity of depolarizing maps is attempted. The geometrical visualisation involves descriptions of the non-Markovianity based on: an unexpected increase in the volume of accessible states of the system signifying the non-Markovian dynamics [22], and employing a trajectory-based visualisation of traced trajectories within the parameter space to capture the dynamical evolution of the depolarizing channel, in turn demonstrating the loss of CP-divisibility and the emergence of non-Markovianity under systematic variations in the system parameters [23].
@@
This work is organized as follows. In Section II, the general depolarizing map including the non-Markovianity parameters is derived. In section III, we describe the dynamical techniques in detecting the non-Markovianity of the depolarizing map by observing the negativity of the eigenvalues of the Choi matrix of the intermediate map, witness for quantum information backflow, and the canonical time-dependent decay rate. Furthermore, we utilize the trace-distance-based distinguishability method to detect the non-Markovianity of the map. In section IV, we describe the geometrical tools in detecting the non-Markovianity through: (a) changing the volume of physical states of the system, (b) employing a trajectory-based visualization within the parameter space to trace trajectories and capture the dynamical evolution of the depolarizing channel. Finally, we conclude in section V.
@@ II General depolarizing map
The depolarizing map is frequently encountered in quantum information theory. It takes an initial state of the system $\rho_{S}(0)$ to a mixture of itself and a maximally mixed state [3, 24, 25, 26, 20]. Mathematically, it can be represented as a quantum channel or quantum operation. In general, the operator sum representation of this map is $\Phi(\rho)=\sum_{i}E_{i}\rho E_{i}^{\dagger}$, where the Kraus operators are:
@@
where $\mathbb{I}$ is the identity operator, and $\sigma_{X}$, $\sigma_{Y}$, $\sigma_{Z}$ are the Pauli matrices. Furthermore, $k$ increases monotonically from 00 (noiseless case) to $1$ (maximal depolarizing) [27]. The most general form of the Kraus operators of the depolarizing map can be written as [28, 20, 27]
@@
To validate the map’s complete positivity, we choose $\alpha\in[0,1]$. For instance, if $\alpha=0$, then Eq. (II) reduces to the Kraus operators of the standard depolarizing map, Eq. (II), and $p$ becomes $k$.
In general, $k$ depends on $p$. Comparing Eq. (II) with Eq. (II), $k(p)$ can be expressed as
@@ III Non-Markovianity using dynamical tools
In this section we characterize the non-Markovianity of the depolarizing map, Eq. (II), by observing the negative eigenvalues of the Choi matrix of the intermediate map and the canonical time-dependent decay rate. In addition to it, the witness operator for non-Markovianity is discussed, and the quantification of resulting memory effects due to information backflow is also presented. Furthermore, we utilize the trace-distance-based distinguishability method to detect this non-Markovianity.
If the dynamical map $\Phi(t,0)$ in Eq. (1), is CP-divisible and also invertible, $i.e.,$ the inverse $\Phi^{-1}(t,0)$ exists for all times from $t=0$ to $t$, then one can write the dynamical map for $t\geq s\geq 0$ into the following composition [7],
$\Phi(t,0)=\Phi(t,s)\circ\Phi(s,0),$ (7)
where the $\circ$” represents the composition between two maps. The map $\Phi(t,s)$ describes the intermediate system-environment interactions and can be defined as
$\Phi(t,s)=\Phi(t,0)\left[\Phi^{-1}(s,0)\right],$ (8)
and needs to be CP. Note that the condition in Eq. (7) is the quantum counterpart to the classic Chapman-Kolgomorov equation [7]. However, if the intermediate map $\Phi(t,s)$ is not CP (NCP) [29], then it is indicative of the usual notion of non-Markovianity [5, 7, 20].
The dynamical map $\Phi(t,0)$ in terms of “divisibility”, represents a memoryless evolution, as a composition of physical maps is indicative of quantum Markovianity. Based on the RHP criterion [7], a dynamics is said to be non-Markovian if it is not CP divisible. This does not require optimization but a normalization is needed to handle the singularity. An alternative method of characterizing non-Markovian dynamics is the BLP method [5, 6]. Here, non-Markovianity is defined in terms of the non monotonic behavior of the trace distance, ||.$||.||_{1},$ and is based on the notion of distinguishability between any two initial states $\rho_{1}$ and $\rho_{2}$. A non monotonic behavior of distinguishability is indicative of a back-flow of information from the environment to the system, a signature of non-Markovianity. This requires optimization of the states and does not require normalization. The information backflow criteria does not distinguish between classical and quantum memory effects. Using [21], quantumness in the observed information backflow can be characterised for the intermediate map.
@@ III.1 The eigenvalues of the Choi matrix of the intermediate dynamics
In general, the dynamical map $\Phi(t,0),t>0$ can be obtained by tomography [3, 7]. For the present case, the general depolarizing map is theoretically known, see Eq. (II). To realize the non-Markovianity of the general depolarizing map Eq. (II), we consider the intermediate map $\Phi(t,s)$, $0\leq s\leq t$. If the intermediate map $\Phi(t,s)$ is not CP, the dynamics will be non-Markovian [7].
In the following, we describe the main steps toward calculating the intermediate map in Eq. (8), $\Phi(t,s)$. The general depolarizing map is written as
$\Phi(t,0)(\rho(0))$ $=E_{\mathbb{I}}\rho(0)E^{\dagger}_{\mathbb{I}}+\sum_{i=X,Y,Z}E_{i}\rho(0)E_{i}^{\dagger}$
@@
Furthermore, we vectorize $\Phi(t,0)$ by writing the resulting matrix as a (column) vector by gathering the columns on top of one another. This process is called “vectorization” [7, 30, 31]. The diagonalized form of the matrix elements of the unital intermediate map obtained using $\Phi(t,0)$, following Eq. (8) is
$\Phi(t,s)=\Phi(t,0)\Phi^{-1}(s,0)=\begin{pmatrix}1&0&0&0\\ 0&\lambda_{1}(p)&0&0\\ 0&0&\lambda_{2}(p)&0\\ 0&0&0&\lambda_{3}(p)\end{pmatrix},$ (10)
@@
Here, $0\leq q\leq p\leq 1$. It is worth mentioning here that when $q=\alpha=0$ in Eq. (11), then $\Phi(t,s)=\Phi(t,0)$, and $\lambda_{1}(p)=\lambda_{2}(p)=\lambda_{3}=(1-p)$. This illustrates the standard depolarizing map, and the map shrinks the Bloch sphere uniformly along $x$, $y$, and $z$ to a radius of $(1-p)$ [26, 24].
To find out whether $\Phi(t,s)$ is NCP or not, we use a Choi matrix constructed by first constructing the matrix $U_{2\leftrightarrows 3}[\Phi(t,s)\otimes\mathbb{I}_{4}]U_{2\leftrightarrows 3}$ where $U_{2\leftrightarrows 3}$ is the commutation (or ”swap“) matrix between the ”second“ and the ”third“ subspace [111]; second, apply $U_{2\leftrightarrows 3}[\Phi(t,s)\otimes\mathbb{I}_{4}]U_{2\leftrightarrows 3}$ on $vec(\ket{\Psi}\bra{\Psi})$, where |$\ket{\Psi}=\frac{1}{\sqrt{2}}\left(\ket{00}+\ket{11}\right)$. $U_{2\leftrightarrows 3}$ is written as follows
@@
We assume that $\phi(t,s)\otimes\mathbb{I}_{4}$ acts in a tensor product of four spaces with the same dimension , $\mathcal{H}_{1}\otimes\mathcal{H}_{2}\otimes\mathcal{H}_{3}\otimes\mathcal{H}_{4}$. Then $U_{2\leftrightarrows 3}$ denotes the permutation matrix interchanging the second and third subspace, $i.e.$ $U_{2\leftrightarrows 3}=\mathbb{I}_{2}\otimes U_{P}\otimes\mathbb{I}_{2}$. Third, we write the result as a matrix, $i.e.$ “devectorize” to construct Choi matrix of the intermediate map as follows
$\chi(\alpha,q,p)=\left[\Phi(t,s)\otimes\mathbb{I}\right]\ket{\Psi}\bra{\Psi}.$ | . (13)
By Choi-Jamiolkowski isomorphism, matrix $\chi(\alpha,q,p)$ is positive if $\Phi(t,s)$ is CP [20, 32]. An NCP $\Phi(t,s)$ is indicative of non-Markovian dynamics. The Choi matrix for the intermediate map, $\chi(\alpha,q,p)$, is found to be
@@
Figure 1: (Color online). The eigenvalues of the Choi matrix (14). Solid, black line($\lambda_{\mathbb{I}}$) and, dashed, black line, ($\lambda_{X,Y,Z}$), both for $\alpha=0.0$. Solid, red line ($\lambda_{\mathbb{I}}$) and, red, dashed line ($\lambda_{\mathbb{I}}$) both for $\alpha=0.7$. The intermediate $p$-range lies between $p=q=0.3$ and $p=p_{max}=1.0$. Here $q<\alpha_{-}$, and crossover at $\alpha_{-}\approx 0.78$
@@
where $i=\{X,Y,Z\}$. In the case of no perturbations, $i.e.,$ $\alpha=0$ and $q=0$, the map is CP-divisible in the range $0\leq p\leq 1$. In this case, the eigenvalues become $\Lambda_{\mathbb{I}}=\left(1-\frac{3}{4}p\right)$, and $\Lambda_{i}=(\frac{p}{4})$. Furthermore, when $p=p_{max}=1.0$, the depolarizing is maximum, indicated by the intersecting points of the black lines in Fig. (1).
On the other hand, introducing a perturbation during the dynamics, initiates some form of memory into the system. These memory effects are controlled by the factor $\alpha$. The eigenvalues of the Choi matrix show crossover for $\alpha\neq 0$ and $q\neq 0$. The timelike $p$ varies from $q$ to $p_{max}=1$. The $p$ satisfying $04p+4\alpha p-3\alpha p^{2}-4=$ 0 gives $\alpha_{\pm}=\frac{2(1+\alpha\pm\sqrt{1-\alpha+\alpha^{2}})}{3\alpha}$, corresponding to points of crossover of eigenvalues. The $\alpha_{+}$ is outside the domain of $p$ and is not considered for study. For $q<\alpha_{-}$, eigenvalues crossover at $\alpha_{-}$. As depicted in Fig. (1), the point $p=\alpha_{-}$ represents maximal depolarizing. For $q>\alpha_{-}$, as illustrated in Fig. (2), $\Lambda_{i}$ become negative in the entire range of $p$ for the intermediate map indicating non-Markovianity. The point $q=\alpha_{-}$ represents a singularity, since $\Lambda_{I}$ and $\Lambda_{X,Y,Z}$ diverge for any $p\in(q,1]$. The negative eigenvalue $\Lambda_{X,Y,Z}$ suggests that the trace norm of Choi matrix, $||\chi(\alpha,q,p)||_{1}$ is greater than 1 in this interval. Since $\Phi(t,s)$ is trace-preserving, $||\chi(\alpha,q,p)||_{1}$ provides a witness of the NCP character of $\Phi(t,s)$ as [7]
$||\chi(\alpha,q,p)||_{1}$ $=||\left[\Phi(t,s)\otimes\mathbb{I}\right]\ket{\Psi}\bra{\Psi}||_{1}$
@@
Integrating $\chi(\alpha,q,p)$ over the evolution time provides a measure of non-Markovianity, the RHP measure [7]. Moreover, NCP intermediate map relates to negative decay rates, suggesting a conceptually equivalent, but quantitatively different and possibly computationally easier method to indicate non-Markovianity, based on the integral of the normalized time dependent decay rate function in the canonical form of the master equation for the negative decay rate(s), the HCLA measure [19], used below.
Figure 2: The eigenvalues of for the Choi matrix (III.1). Sold line ($\lambda_{\mathbb{I}}$) and, dashed line ($\lambda_{X,Y,Z}$) , for $\alpha=0.7$ (non-Markovian). The intermediate $p$-range lies between $q=0.8$ and $p_{max}=1.0$. The whole range of $p$ corresponds to an NCP map, demonstrating the non-Markovianity of the channel. Here $q>\alpha_{-}$ as for $\alpha=0.7$ corresonding $\alpha_{-}\approx 0.78$.

### v2

@@ Abstract
The non-Markovian depolarizing channel is explored from the perspective of understanding its non-Markovian behavior as well as the occurrence of singularities. The study brings together the various ways to identify and quantify non-Markovianity. This includes dynamical techniques such as quantum information backflow witness, Breuer-Laine-Piilo, Rivas-Huelga-Plenio and Hall-Cresser-Li-Andersson measures. In addition, geometrical visualization of non-Markovian effects is presented using the variation in the volume of accessible states during dynamical evolution. Further, a trajectory-based visualization of the dynamical map within the parameter space is presented. The trajectories traced during evolution demonstrate the loss of CP divisibility and the emergence of non-Markovianity under systematic variations of the system parameters. The effects of increasing system dimensions and qubit numbers on singularity and non-Markovianity are presented, with an extension of characterization techniques to higher-dimensional systems.
@@ I Introduction
The dynamical evolution of the open quantum system [1, 2], a composite system consisting of a smaller subsystem of interest and an interacting larger subsystem called environment, is of interest to the field of quantum information [3, 4] due to arising memory effects and non-Markovianity [5, 6, 7, 8, 9, 10, 11, 12]. The concept of the completely positive trace preserving (CPTP) map is used to study such evolution, which maps the systems initial state $\rho_{S}(0)$ to the evolved final state $\rho_{S}(t)$ [3, 13]
@@
Alternatively, CPTP maps are called quantum channels and can be described using operator sum representation (Kraus formalism) [14, 15, 16, 17]. The mapped evolution by the CPTP map can be classified as Markovian or non-Markovian [18]. In Markovian evolution, changes in the system state depend solely on its current state, with no influence from past events. Conversely, in non-Markovian evolution, this independence is breached, indicating that the environment retains a memory of the system history, affecting its future evolution.
To identify non-Markovianity, traditionally, one resorts to the loss of CP divisibility [7] or nonmonotonic increase in the distinguishability of a pair of states during evolution [6]. Even CP-divisible processes can exhibit non-Markovian effects [9, 10]. The increase in distinguishability is indicative of information backflow from environment to system, indicating non-Markovianity. The Breuer-Laine-Piilo (BLP) measure [5] uses distinguishability to quantify non-Markovianity. However, the BLP measure is not the only method to detect non-Markovianity. There are other witnesses, such as the Rivas-Huelga-Plenio (RHP) measure [19, 7] and Hall-Cresser-Li-Anderson (HCLA) measure [20], as discussed below. The divisibility of a dynamical map can be used to characterize non-Markovianity. The dynamical map is said to be CP-indivisible, if the intermediate map (propagator), is not completely positive (NCP), possessing at least one negative eigenvalue for the corresponding Choi matrix. The intermediate NCP map is indicative of system-environment correlation. The RHP measure [19, 7] uses the previously mentioned criteria to quantify non-Markovianity. Similarly, the HCLA measure [20] quantifies non-Markovian effects but relies on negative time-dependent decay rates, which are indicative of non-Markovianity. Both measures are equivalent to each other and provide similar insights into non-Markovian processes.
@@
The study of non-Markovianity in quantum channels is necessary for understanding the quantum information processes. Non-Markovian dynamics can significantly affect the quantum systems ways of information preservation and manipulation. By studying non-Markovian effects, better protocols can be developed to mitigate errors [23, 24] in quantum systems, enhancing the reliability and capabilities of quantum computations and communications.
This paper adds to this growing body of research, addressing current gaps in understanding. Moreover, as quantum technology [25, 26] rapidly advances, the necessity of comprehending and understanding non-Markovian effects increases.
Exploring non-Markovianity in open quantum systems is crucial in today’s world, as modern technologies increasingly rely on non-Markovian processes. The importance of this study is driven by both foundational and practical considerations. Fundamentally, exploring the non-Markovian nature of system-environment interactions is essential for accurately describing a variety of real-world systems. Examples include quantum biological systems [27], complex quantum networks [28], solid-state devices such as superconductor quantum interference devices (SQUIDs) and Josephson junctions [29], and ultracold gases [30], all of which exhibit non-Markovian behavior. Incorporating non-Markovian effects into the analysis of these physical systems allows for more precise and reliable predictions.
In engineering and control systems [31], incorporating non-Markovian dynamics can significantly enhance the performance and stability of control mechanisms. Similarly, in material science [32], non-Markovian models play a crucial role in understanding and designing materials with tailored properties, as memory effects can influence their mechanical, thermal, and electrical characteristics, facilitating advancements in material design. By addressing both theoretical and practical aspects, the study of non-Markovian dynamics effectively bridges the gap between academic research and real-world applications, thereby advancing scientific understanding and fostering technological innovation.
The depolarizing channel is a fundamental noise model in quantum information theory, representing a type of noise that occurs in quantum systems. Analyzing its non-Markovianity can give insights into error correction [33, 3, 34], quantum communication [35, 36], and quantum computing [33, 3].
The depolarizing channel can be produced by several types of environments, generally characterized by their randomizing effects on the quantum state. For example, in environments where the system is subjected to random fluctuations or noise, such as thermal noise, background radiation, or uncontrolled electromagnetic fields, the qubit can be depolarized [37, 38, 39, 40]. These fluctuations cause random rotations in the state of the qubit, leading to a mixture of all possible states, which is characteristic of the depolarizing channel. Furthermore, in optical quantum systems, depolarization can occur due to scattering processes. For example, when photons interact with particles in a medium, they can scatter in a way that randomizes their polarization, leading to a depolarized state [41, 42]. Thus, any environment that induces random, unbiased noise or interactions on a quantum system can give rise to a depolarizing channel. This randomness leads to the equal probability of all possible outcomes, which is the defining feature of depolarization.
Here, we try to establish a foundational understanding of non-Markovianity in the depolarizing channel. The simplicity of a one-qubit system allows for clearer interpretation and validation of theoretical models. Additionally, an extension of depolarizing noise to $N$-level qudit systems [43, 44, 45, 46] and multiqubit systems is made, allowing for the study of non-Markovianity in higher-dimensional systems.
@@
Our goal in this article is to answer the following questions: (1) Under what conditions does the depolarizing map become non-Markovian? (2) Can we quantify the amount of this non-Markovianity? and (3) What will be the effect of the increased levels of system or number of qubits on non-Markovianity? In order to answer these questions, both dynamical (based on CP divisibility, distinguishability, negative time-dependent decay rates, and quantum information backflow witness) and geometrical techniques are employed, and a comprehensive understanding of the non-Markovianity of depolarizing maps is attempted. The geometrical visualization involves descriptions of the non-Markovianity based on: an unexpected increase in the volume of accessible states of the system signifying the non-Markovian dynamics [47], and employing a trajectory-based visualization of traced trajectories within the parameter space to capture the dynamical evolution of the depolarizing channel, in turn demonstrating the loss of CP divisibility and the emergence of non-Markovianity under systematic variations in the system parameters [48].
@@
This work is organized as follows. In Sec. II, the general depolarizing map including the non-Markovianity parameters is derived. In Sec. III, we describe the dynamical techniques in detecting the non-Markovianity of the depolarizing map by observing the negativity of the eigenvalues of the Choi matrix of the intermediate map, witness for quantum information backflow, and the canonical time-dependent decay rate. Furthermore, we utilize the trace-distance-based distinguishability method to detect the non-Markovianity of the map. In Sec. IV, we describe the geometrical tools in detecting the non-Markovianity through: (a) changing the volume of physical states of the system, and (b) employing a trajectory-based visualization within the parameter space to trace trajectories and capture the dynamical evolution of the depolarizing channel. In addition, we also characterize non-Markovianity in a qutrit system, thereby paving the way for understanding non-Markovian open quantum systems in higher dimensions, in Sec. V. The non-Markovian behavior in multiqubit systems is presented in Sec. VI.Finally, we conclude in Sec. VII.
@@ II General depolarizing map
The depolarizing map can be defined as a quantum operation that transforms the density matrix of the initial state of the system into a convex combination of the original density matrix with probability of $1-k$ and the maximally mixed state with probability $k$ [3, 21]. Mathematically, it is expressed as $\Phi(\rho)=\sum_{i=\mathbb{I},X,Y,Z}E_{i}\rho E_{i}^{\dagger}$, where
@@
where $\mathbb{I}$ is the identity operator; $\sigma_{X}$, $\sigma_{Y}$, $\sigma_{Z}$ are the Pauli matrices; and $\sum_{i}E_{i}^{\dagger}E_{i}=\mathbb{I}$. Furthermore, $k$ increases monotonically from 00 (noiseless case) to $1$ (maximal depolarizing) [49]. The most general form of the Kraus operators of the depolarizing map can be written as [37, 21, 49]
@@
To validate the map’s complete positivity, we choose $\alpha\in[0,1]$. For instance, if $\alpha=0$, then Eq. (II) reduces to the Kraus operators of the standard depolarizing map, Eq. (II), and $p$ becomes $k$. In general, $k$ depends on $p$. Comparing Eq. (II) with Eq. (II), $k(p)$ can be expressed as
@@ III Non-Markovianity using dynamical tools
In this section, we characterize the non-Markovianity of the depolarizing map, given by Eq. (II), by observing the negative eigenvalues of the Choi matrix of the intermediate map and the canonical time-dependent decay rate. In addition, the witness operator for non-Markovianity is discussed, and the quantification of resulting memory effects due to information backflow is also presented. Furthermore, we utilize the trace-distance-based distinguishability method to detect this non-Markovianity.
If the dynamical map $\Phi(t,0)$ in Eq. (1) is CP divisible, then it can be represented as a series of propagators, described by CP maps $\Phi(t,s)$, $0\leq s\leq t$,
$\Phi(t,0)=\Phi(t,s)\circ\Phi(s,0).$ (7)
Furthermore, for invertible $\Phi(t,0)$ the propagator $\Phi(t,s)$ is well defined as [7, 50]
$\Phi(t,s)=\Phi(t,0)\left[\Phi^{-1}(s,0)\right].$ (8)
Note that the condition in Eq. (7) is the quantum counterpart to the classic Chapman-Kolgomorov equation [7]. However, if the intermediate map $\Phi(t,s)$ is not CP (NCP) [51], then it is indicative of the usual notion of non-Markovianity [5, 7, 21].
The dynamical map $\Phi(t,0)$ in terms of “divisibility” represents a memoryless evolution, as a composition of physical maps is indicative of quantum Markovianity. Based on the RHP criterion [7], a dynamics is said to be non-Markovian if it is not CP divisible. This does not require optimization but a normalization is needed to handle the singularity. An alternative method of characterizing non-Markovian dynamics is the BLP method [5, 6]. Here, non-Markovianity is defined in terms of the non monotonic behavior of the trace distance, ||.$||.||_{1},$ and is based on the notion of distinguishability between any two initial states $\rho_{1}$ and $\rho_{2}$. A non monotonic behavior of distinguishability is indicative of a backflow of information from the environment to the system, a signature of non-Markovianity. This requires optimization of the states and does not require normalization. The information backflow criteria does not distinguish between classical and quantum memory effects. Using [22], quantumness in the observed information backflow can be characterized for the intermediate map.
@@ III.1 The eigenvalues of the Choi matrix of the intermediate dynamics
Although the dynamical map $\Phi(t,0),t>0$, can be obtained through tomography [3, 7], the exponential scaling of this process makes it costly. However, detecting non-Markovianity does not necessarily require such procedures. By establishing accurate lower and upper bounds for properties such as trace distance [6] or entanglement [7] using simpler measurements, we can detect nonmonotonic behavior without relying on expensive tomography. Also, advancements in the field of open quantum systems report possible characterization of dynamical maps even for higher dimensions [52, 53, 54]. For the present case, the general depolarizing map is theoretically known; see Eq. (II). To realize the non-Markovianity of the general depolarizing map given by Eq. (II), we consider the intermediate map $\Phi(p,q)$, $0\leq q\leq p$. If the intermediate map $\Phi(p,q)$ is not CP, the dynamics will be non-Markovian [7].
In the following, we describe the main steps toward calculating the intermediate map in Eq. (8), $\Phi(p,q)$. The general depolarizing map is written as
$\Phi(p,0)(\rho(0))$ $=E_{\mathbb{I}}\rho(0)E^{\dagger}_{\mathbb{I}}+\sum_{i=X,Y,Z}E_{i}\rho(0)E_{i}^{\dagger}$
@@
Furthermore, we vectorize $\Phi(p,0)$ by writing the resulting matrix as a (column) vector by gathering the columns on top of one another. This process is called “vectorization” [7, 55, 56].
The diagonalized form of the matrix elements of the unital intermediate map obtained using $\Phi(p,0)$, following Eq. (8) is
$\Phi(p,q)=\Phi(p,0)\Phi^{-1}(q,0)=\begin{pmatrix}1&0&0&0\\ 0&\lambda_{1}(p)&0&0\\ 0&0&\lambda_{2}(p)&0\\ 0&0&0&\lambda_{3}(p)\end{pmatrix},$ (10)
@@
Here, $0\leq q\leq p\leq 1$, where $p$ and $q$ are regarded as ’timelike’ variables, in place of using time $s$ and time $t$. It is worth mentioning here that when $q=\alpha=0$ in Eq. (11), then $\Phi(p,q)=\Phi(p,0)$, and $\lambda_{1}(p)=\lambda_{2}(p)=\lambda_{3}=(1-p)$. This illustrates the standard depolarizing map, and the map shrinks the Bloch sphere uniformly along $x$, $y$, and $z$ to a radius of $(1-p)$ [57, 33].
To find out whether or not $\Phi(p,q)$ is NCP, we use a Choi matrix constructed by first constructing the matrix $U_{2\leftrightarrows 3}[\Phi(p,q)\otimes\mathbb{I}_{4}]U_{2\leftrightarrows 3}$ where $U_{2\leftrightarrows 3}$ is the commutation (or “swap”) matrix [58, 56] between the “second” and the “third” subspace [7]; second, apply $U_{2\leftrightarrows 3}[\Phi(t,s)\otimes\mathbb{I}_{4}]U_{2\leftrightarrows 3}$ on $\rm{vec}(\ket{\Psi}\bra{\Psi})$ (vectorization of |$\ket{\Psi}\bra{\Psi}$ |), where |$\ket{\Psi}=\frac{1}{\sqrt{2}}\left(\ket{00}+\ket{11}\right)$. $U_{2\leftrightarrows 3}$ is written as follows
@@
We assume that $\phi(p,q)\otimes\mathbb{I}_{4}$ acts on the tensor product of four spaces with the same dimension, $\mathcal{H}_{1}\otimes\mathcal{H}_{2}\otimes\mathcal{H}_{3}\otimes\mathcal{H}_{4}$. Then operator $U_{2\leftrightarrows 3}$ is the permutation matrix causing the interchange of the second and third subspace. Third, we write the result as a matrix, $i.e.$ “devectorize” to construct the Choi matrix of the intermediate map as follows
$\chi(\alpha,q,p)=\left[\Phi(p,q)\otimes\mathbb{I}\right]\ket{\Psi}\bra{\Psi}.$ | . (13)
The generalization of this process for the $N$-level system is presented in the Appendix A. By Choi-Jamiolkowski isomorphism, matrix $\chi(\alpha,q,p)$ is positive if and only if $\Phi(p,q)$ is CP [21, 59]. A nonpositive semidefinite Choi matrix of propagator indicates non-Markovianity. The Choi matrix for the intermediate map, $\chi(\alpha,q,p)$, is found to be
@@
Figure 1: (Color online). The eigenvalues of the Choi matrix (14). Solid-black line ($\Lambda_{\mathbb{I}}$) and dashed black line ($\Lambda_{X,Y,Z}$), both for $\alpha=0.0$. Solid red line ($\Lambda_{\mathbb{I}}$) and red dashed line ($\Lambda_{\mathbb{I}}$) both for $\alpha=0.7$. The intermediate $p$ range lies between $p=q=0.3$ and $p=p_{max}=1.0$. Here,$q<\alpha_{-}$, and crossover at $\alpha_{-}\approx 0.78$
@@
where $i=\{X,Y,Z\}$. In the case of no perturbations, $i.e.,$ $\alpha=0$, and $q=0$, the map is CP-divisible as it reduces to a simple case of a Markovian depolarizing channel. In this case, the eigenvalues become $\Lambda_{\mathbb{I}}=\left(1-\frac{3}{4}p\right)$ and $\Lambda_{i}=(\frac{p}{4})$. The same is true for $\alpha\neq 0$ and $q=0$, just presenting the perturbed depolarizing channel. Furthermore, when $p=p_{max}=1.0$, the depolarizing is maximum, indicated by the intersecting points of the black lines in Fig. 1.
On the other hand, introducing a perturbation during the dynamics, initiates some form of memory into the system. These memory effects are controlled by the factor $\alpha$. The eigenvalues of the Choi matrix show crossover for $\alpha\neq 0$ and $q\neq 0$. The timelike $p$ varies from $q$ to $p_{max}=1$. The $p$ satisfying $04p+4\alpha p-3\alpha p^{2}-4=$ 0 gives $\alpha_{\pm}=\frac{2(1+\alpha\pm\sqrt{1-\alpha+\alpha^{2}})}{3\alpha}$, corresponding to points of crossover of eigenvalues. The $\alpha_{+}$ is outside the domain of $p$ and is not considered for study. For $q<\alpha_{-}$, eigenvalues crossover at $\alpha_{-}$. As depicted in Fig. 1, the point $p=\alpha_{-}$ represents maximal depolarizing. For $q>\alpha_{-}$, as illustrated in Fig. 2, $\Lambda_{i}$ become negative in the entire range of $p$ for the intermediate map indicating non-Markovianity. The point $q=\alpha_{-}$ represents a singularity as the propagator or intermediate map is undefined at this point; the $\Lambda_{\mathbb{I}}$ and $\Lambda_{X,Y,Z}$ diverge for any $p\in(q,1]$. The negative eigenvalue $\Lambda_{X,Y,Z}$ suggests that the trace norm of the Choi matrix, $||\chi(\alpha,q,p)||_{1}$ is greater than 1 in this interval. Since $\Phi(p,q)$ is trace preserving, $||\chi(\alpha,q,p)||_{1}$ provides a witness of the NCP character of $\Phi(p,q)$ as [7]
$||\chi(\alpha,q,p)||_{1}$ $=||\left[\Phi(p,q)\otimes\mathbb{I}\right]\ket{\Psi}\bra{\Psi}||_{1}$
@@
Integrating $\chi(\alpha,q,p)$ over the evolution time provides a measure of non-Markovianity, $i.e.,$ the RHP measure [7]. This quantifies non-Markovian effects in the CP-indivisible regime. Moreover, the NCP intermediate map relates to negative decay rates, suggesting a conceptually equivalent, but quantitatively different and possibly computationally easier method to indicate non-Markovianity, based on the integral of the normalized time dependent decay rate function in the canonical form of the master equation for the negative decay rate(s), $i.e.,$ the HCLA measure [20], used below.
Figure 2: The eigenvalues of the Choi matrix (III.1). Sold line ($\Lambda_{\mathbb{I}}$) and, dashed line ($\Lambda_{X,Y,Z}$) , for $\alpha=0.7$ (non-Markovian). The intermediate $p$ range lies between $q=0.8$ and $p_{max}=1.0$. The whole range of $p$ corresponds to an NCP map, demonstrating the non-Markovianity of the channel. Here $q>\alpha_{-}$ as for $\alpha=0.7$ corresponding with $\alpha_{-}\approx 0.78$.